The concentrated sulfuric acid we use in the laboratory is 98.0% sulfuric acid by mass. Calculate the molarity of the acid solution. The density of the solution is 1.33 g/mL.


1
Expert's answer
2020-12-10T07:52:08-0500

d = 1.33 g/mL

M(H2SO4) = 98.08 g/mol

"n(total) = \frac{d}{M} = \frac{1.33}{98.08} = 0.0135 \;mol"

98.0% means 98.0 g in 100 mL of solution

mass percent "= \frac{n(acid)}{n(total)} \times 100"

"98.0 = \frac{n(acid)}{0.0135} \times 100"

n(acid) = 0.01323 mol/mL

Answer: 13.23 M